# A Survey of Artificial Intelligence in Gait-Based Neurodegenerative Disease Diagnosis

### v1

@@ Abstract.
Recent years have witnessed an increasing global population affected by neurodegenerative diseases (NDs), which traditionally require extensive healthcare resources and human effort for medical diagnosis and monitoring. As a crucial disease-related motor symptom, human gait can be exploited to characterize different NDs. The current advances in artificial intelligence (AI) models enable automatic gait analysis for NDs identification and classification, opening a new avenue to facilitate faster and more cost-effective diagnosis of NDs. In this paper, we provide a comprehensive survey on recent progress of machine learning and deep learning based AI techniques applied to diagnosis of five typical NDs through gait. We provide an overview of the process of AI-assisted NDs diagnosis, and present a systematic taxonomy of existing gait data and AI models. Through an extensive review and analysis of 164 studies, we identify and discuss the challenges, potential solutions, and future directions in this field. Finally, we envision the prospective utilization of 3D skeleton data for human gait representation and the development of more efficient AI models for NDs diagnosis. We provide a public resource repository to track and facilitate developments in this emerging field: https://github.com/Kali-Hac/AI4NDD-Survey.
@@
Artificial intelligence, Neurodegenerative diseases, Gait, Disease diagnosis
journal: CSURjournalvolume: 0journalnumber: 0article: 0publicationmonth: 0journalyear: 0ccs: Computing methodologies Artificial intelligenceccs: Applied computing Health informaticsccs: General and reference Surveys and overviews
1. Introduction
@@
Neurodegenerative Diseases (NDs) such as Alzheimer’s disease (AD) and Parkinson’s disease (PD) are among the most widespread and devastating disorders affecting millions of people worldwide (scheltens2021alzheimer, ). According to the Alzheimer’s Disease Association (AD_statistic, ) and the Parkinson’s Foundation Report (PD_statistic, ) in 2023, nearly 6.2 million people and 1 million people are diagnosed with AD and PD respectively in the United States. While in China, there are over 9.8 million AD patients (ren2022china, ), and it is estimated that Chinese PD patients will increase to 4.9 million by 2030, accounting for a half of the worldwide PD population (dorsey2007projected, ). From 1990 to 2019, the effected global population of AD, PD, and other neurodegenerative diseases ($e.g.$, amyotrophic lateral sclerosis) has significantly increased (illustrated in Fig. 2), leading to a growing burden on global healthcare system. On the one hand, these diseases could result in pathological gaits and chronic pain in body joints, tissues, and nerves (de2016pain, ; negre2008chronic, ), which severely reduce the flexibility, stability, mobility, and other functional capabilities of a human body. On the other hand, they also impose large psychological and financial burdens to patients and their family (beiske2009pain, ; ovaska2021literature, ; averill2007psychological, ). The treatment cost for these diseases has reached approximately USD 130 billion per year and further rise has been estimated (adejare2016drug, ; zahra2020global, ). Therefore, it is essential to conduct an earlier and more reliable diagnosis of these diseases, so as to provide a timely proper treatment to mitigate the burden of societal and healthcare resources.
@@
Figure 1. Worldwide effects (1990 versus 2019) of two most representative neurodegenerative diseases (AD and PD), measured in “Disability-Adjusted Life Years (DALYs)” (ding2022global, ).
@@
Figure 2. The number of studies (from 2012 to 2022) relevant to gait-based neurodegenerative disease diagnosis using AI.
@@
As one of the most essential motor symptoms associated with pathology, locomotion anomalies, especially observed in human gait, can reflect the incidence and progression of different NDs (kirchner2014detrended, ; sejdic2013comprehensive, ; tang2013locomotion, ). Taking AD, PD and amyotrophic lateral sclerosis (ALS) as an example, these diseases often cause patients to exhibit marked alterations and abnormalities in their walking patterns (kobsar2017wearable, ; rutherford2018knee, ; spasojevic2015vision, ; armand2016gait, ; sanders2010gait, ; allet2008gait, ; myers2016gait, ; myers2013vascular, ; wurdeman2012patients, ; szymczak2018gait, ). In general, compared with a normal gait (see Fig. 4), an impaired gait may display changes in speed, cadence, and limb moments within a gait cycle (favre2014age, ; creaby2012gait, ; spasojevic2015vision, ). As shown in Fig. 4, the gait with a widened base, unsteadiness and irregularity of steps, and lateral veering (termed Parkinsonian gait) could suggest PD (spasojevic2015vision, ; ebersbach2013clinical, ). In the case of advanced ALS, the patients may possess foot drop, where one foot flops down when lifting the leg, necessitating a higher lifting of the knee to prevent the toe from dragging on the ground (termed neuropathic gait) (hausdorff2000dynamic, ; isaacs2007amyotrophic, ; wijesekera2009amyotrophic, ).
@@
In recent years, driven by the widespread availability of body positioning technologies and economical portable devices to monitor human gait, gait-based NDs diagnosis ($i.e.,$ NDs classification, identification or prediction) has propelled a surge of attention across academic, industrial, and medical communities (aich2020design, ; khera2022age, ; fraiwan2021computer, ; kour2023sensor, ; di2020gait, ; pardoel2019wearable, ; gupta2023new, ; vienne2017inertial, ; das2022recent, ; salchow2022emerging, ; loh2021application, ; ayaz2023automated, ; zolfaghari2023sensor, ). To perform NDs diagnosis from diverse gait data, the Artificial Intelligence (AI) technologies, especially Machine Learning (ML) and Deep Learning (DL) models, have been extensively explored and applied (kour2022vision, ; khera2022age, ; fraiwan2021computer, ; kour2019computer, ; kour2023sensor, ; di2020gait, ). A common practice is to extract gait parameters ($e.g.,$ stride length, stance duration, gait cycle time) from sensor-based data such as vertical ground reaction force (vGRF) or vision-based data such as gait images, and leverage ML algorithms ($e.g.,$ support vector machine (SVM)) or deep neural networks to automatically recognize gait patterns associated with NDs to aid in their diagnosis (mannini2016machine, ; wahid2015classification, ; abdulhay2018gait, ). These methods not only circumvent manual observation and assessment of NDs patterns in traditional diagnosis approaches, but also show immense potential in assisting healthcare professionals to diagnose and treat these diseases more efficiently (kour2023sensor, ; di2020gait, ). In this paper, we systematically review recent advancements of utilizing AI to help diagnose NDs from human gait. In particular, our work focuses on the diagnosis of five most prevalent and typical NDs (PD, AD, ALS, Huntington’s Disease (HD), Multiple Sclerosis (MS)) and all existing AI models that can help diagnose these diseases ($i.e.,$ identify, classify or predict NDs) from human gait data. A total of 2069 potentially relevant articles, including 14 survey papers, are found from three databases (PubMed, Web of Science (WoS), Google Scholar). We first screen these articles by titles and abstracts based on the pre-defined inclusion criteria, and then conduct a full-text assessment to exclude irrelevant papers. Finally, 164 of 2069 papers are selected to be included in our survey.
@@
Figure 3. Different phases of normal human gait with coordinated movements of legs and arms. Right: Examples for hemiplegic gait.
@@
Figure 4. Examples for hemiplegic gait (left), Parkinsonian gait (middle), and neuropathic gait (right) (https://doi.org/10.1002/dac.4348, ).
1.1. Objectives and Contributions
@@
The objectives of this survey can be encapsulated in the following six key questions (Q): (Q1) How is human gait associated with NDs? (Q2) What is the process of AI-assisted NDs diagnosis from gait data? (Q3) What are various gait data types, gait collecting approaches, and AI model types for NDs diagnosis? (Q4) What are state-of-the-art AI models and their performance results ($e.g.,$ diagnostic accuracy) in gait-based NDs diagnosis? (Q5) What are the main challenges and promising future directions of this field? (Q6) Is it feasible to exploit 3D skeleton data to represent human gait and build AI models for NDs diagnosis?
@@ 2. Preliminaries
2.1. Basic Concepts
  1. (1)
    Gait: It is defined as a mechanism of locomotion, which contains rhythmic and coordinated movements of different limbs to enable the forward progression of the body (sethi2022comprehensive, ; owusu2007ai, ) (as shown in Fig. 4). The gait of a person typically consists of the motion of lower extremities and their correlated posture in upper limbs or torso, which normally requires a synthesized coordination of nerves, skeletons, and muscles (Nutt268, ). Owing to this reason, the abnormality or impairment of gait can serve as an effective indicator of potential body pains or diseases caused by physical injury, aging or related disorders (kour2021survey, ; loh2021application, ).
  2. (2)
@@ 2.1. Basic Concepts
  5. (5)
    Disease Management: General disease management methods consists of therapies using a combination of medications, physical therapies, psychological therapies, interventional procedures or surgeries. For NDs, the management is often disease-specific with focus on either the disease pathogenesis or symptoms experienced (lamptey2022review, ).
2.2. Pathological Gaits
2.2.1. Neurodegenerative Disease Related Gaits
@@
2.2.2. Other Disease-Related Gaits
@@
Figure 5. Overview for the process of AI-assisted gait-based neurodegenerative disease diagnosis.
Figure 6. The taxonomy of gait data types used for AI-assisted NDs diagnosis.
Figure 7. Overview of AI model learning workflow.
2.3. Overview of AI-assisted Gait-Based Neurodegenerative Disease Diagnosis
@@
It is noteworthy that exploiting AI models for gait analyses and classification ($i.e.$, Step 2) is the main part of the gait-based NDs diagnosis process. It includes the workflow of designing, training, and validating AI models using pre-collected gait data. As shown in Fig. 7, the AI learning workflow can be divided into (a) raw gait data collection, (b) data preprocessing or/and augmentation, (c) model design or selection, (d) model training and validation, (e) model testing and deployment, and (f) iterative improvement ($e.g.,$ model fine-tuning using more data). For conventional ML models, feature engineering is usually required after step (c) to manually extract discriminative feature such as gait descriptors. For DL models, the design of models ($e.g.$, neural network architectures) (step (c)) and the sufficiency of training data (step (a) and (b)) are often the two most essential determinants of model performance. Since more high-quality training data can better guide the model optimization, extend the learnable feature space, and improve the generalization ability ($i.e.,$ higher performance on new testing data) of the model, it is feasible to collect more gait data while using effective data augmentation strategies ($i.e.$, generate more augmented training samples) to iteratively boost the model performance (detailed in Sec. 5.1).
2.4. Gait Data Taxonomy
@@
  • Sensor Modality: Advancement in sensor technologies allows its application in various tasks such as gait analysis, action recognition, etc. The sensors explored among the collected papers offer various advantages and could be classified into these four types, including force, inertial, hybrid, and other less prevalent sensors (alam2017vertical, ; hu2022machine, ; kaur2020predicting, ; perumal2016gait, ; das2022recent, ).
    1. (1)
@@
      Force Sensor (FS): There are mainly two types of force or pressure capturing approaches. The first is to use non-wearable floor-mounted devices ($e.g$, force mats, treadmill) to measure vertical Ground Reaction Force (vGRF), moments, or plantar pressure profiles (das2022recent, ). In a laboratory setting, these stationary floor-mounted devices are sufficient for precise data collection. However, since those devices are often bulky, operating-costly, and expensive, it’s not often clinically applicable. Therefore, researchers develop some small wearable devices ($e.g.$, insole) whose application in the field has became a popular choice and risen significantly in the past decade, due to its portable features and cheap installment, at the cost of less accuracy and reliability compare with the floor-mounted ones (howell2012insole, ; howell2013kinetic, ).
    2. (2)
@@
      Inertial Sensor: Also termed as Inertial Measurement Unit (IMU), is often attached to limb segments ($e.g$, foot) (carvajal2022classification, ) or embedded in wearable devices ($e.g.$, smartwatch) to capture dynamic motion ($e.g.$, 3D-acceleration) of the subject (arora2014high, ; klucken2013unbiased, ). These sensors or devices record the repeating pattern of gait signals generated by the locomotion of the subjects. These signals will then be interpreted with various algorithms for the intended tasks, including gait event detection, kinetic or kinematic parameters estimation, and gait classification.
    3. (3)
@@
      Multi-Sensor: These sensors typically combine various types of sensors for a more comprehensive data collection. Among the surveyed papers, studies such as (perez2020identification, ; mileti2018measuring, ; li2022abnormal, ) utilize both IMU sensors and Force Sensitive Resistors (FSRs), while (guo2019method, ) employs Electroencephalogram (EEG) sensors and FSRs. These combinations harvest the strengths of each sensor type hence help enhancing the reliability and accuracy of gait event detection and classification.
    4. (4)
@@
      Other Sensors: Sensors included in this category are less prevalent and exploited among collected papers. However, these sensors still play an important role in providing a more full-sided understanding of the human gait. For instance, EEG sensor is used to measure the brain electrical activity signals data which can be used to analyze brain activity patterns that may correlate with specific gait abnormalities (zhao2023wearable, ; guo2019method, ; you2020alzheimer, ). Electromyography (EMG) sensor is another useful sensor which is used to measure the electrical activity produced by muscles whose activities are controlled by the motor neurons (kugler2013automatic, ). Since NDs patients suffer from the degeneration of nerve cells, they manifest significantly different muscle electrical pattern from healthy subjects (kour2023sensor, ). Therefore, their EMG data can be utilized to aid NDs diagnosis (kugler2013automatic, ).
@@
  • Vision Modality: The vision modality employs optoelectronic Motion capture (Mocap) systems, including various types of cameras ($e.g.$, Infrared (IR) Camera (park2021classification, )), to capture the locomotion of subjects during walking. It’s used to overcome the limitations of the traditional clinical examination which purely depends on the physicians’ naked eyes. Depends on the precision requirement of the video data, the vision modality could be further divided into two sub-categories, namely marker-based and marker-less systems (kour2019computer, ).
    1. (1)
@@
      Marker-based: This approach is developed at an early stage when the cameras are not advanced enough for precise motion capture. Therefore, these Mocap systems ($e.g.$, Vicon) use reflective markers to help capture targets’ movement in the images more precisely. The advantages of using the marker-based approach are that they’re highly accurate and provides exact location of the body landmark. These advantages hence make this modality a gold standard in the field. However, this method requires attaching markers to the body, leading to additional expenses and the potential of alter the gait (filtjens2021modelling, ; aich2018machine, ; ferreira2022machine, ).
    2. (2)
@@
  • Sensor+Vision Modality: This modality could also be referred to as multi-modal modality. In comparison with single-modal approaches, although the latter ones have their own merits and show some promising results, this combined approach unleashes a bigger potential, such as providing richer data and more comprehensive features, capturing more subtle changes in gait, improving recognition performance (you2020alzheimer, ; zhao2021multimodal, ; chatzaki2022can, ). In particular, the application of Microsoft Kinect system has gain a tremendous amount of attention in the field due to its utilization of both depth sensor and RGB camera. This integration enables a more accurate estimation of the gait parameters without sacrificing the advantages mentioned in the marker-less vision modality. More specifically, using conventional cameras to estimate skeleton coordinates produces coarse results. However, when employing the depth sensor, the spatiotemporal features can be precisely derived from the skeleton joints’ coordinates obtained using it. The precision and efficacy of using Kinect has been verified in many studies which compare the results with the gold standard (marker-based vision) approach. Nevertheless, one limitation of current Kinect system is that its precision reduces as the subject walks further away from the camera, resulting in a width and length-limited data collection space (buongiorno2019low, ; tupa2015motion, ; reyes2019lstm, ; munoz2022machine, ). Besides the RGB-D camera approach, there are also other combinations of these two modalities. For example, Tahir et al. fused a marker-based IR camera and a force-sensitive platform (tahir2012parkinson, ), while Chatzaki et al. utilized a conventional camera, IMU, and a force platform (chatzaki2022can, ). Additionally, Zhao et al. merge some vGRF and IMU sensors with Kinect (zhao2021multimodal, ).
  • Other Modalities: Trabassi et al. use Electronic Medical Record (EMR) which contains various patient-related features and medical history data that are indicative of early PD symptoms or risk factors as well as the first diagnosis of gait or tremor disorders to create prediction models (yuan2021accelerating, ).
@@
Figure 8. Statistics of the included papers with different (a) AI model types, (b) NDs types, (c) experimental sample sizes, (d) gait data types, (e) sensor types, and (f) countries. We show the lowest and top three quantities in (f). Papers of different years are shown in Fig. 2. Detailed analyses are provided in Sec. 3.2.
2.5. AI Model Taxonomy

### v2

@@ Abstract
Recent years have witnessed an increasing global population affected by neurodegenerative diseases (NDs), which traditionally require extensive healthcare resources and human effort for medical diagnosis and monitoring. As a crucial disease-related motor symptom, human gait can be exploited to characterize different NDs. The current advances in artificial intelligence (AI) models enable automatic gait analysis for NDs identification and classification, opening a new avenue to facilitate faster and more cost-effective diagnosis of NDs. In this paper, we provide a comprehensive survey on recent progress of machine learning and deep learning based AI techniques applied to diagnosis of five typical NDs through gait. We provide an overview of the process of AI-assisted NDs diagnosis, and present a systematic taxonomy of existing gait data and AI models. Meanwhile, a novel quality evaluation criterion is proposed to quantitatively assess the quality of existing studies. Through an extensive review and analysis of 164 studies, we identify and discuss the challenges, potential solutions, and future directions in this field. Finally, we envision the prospective utilization of 3D skeleton data for human gait representation and the development of more efficient AI models for NDs diagnosis [^1] We provide a public resource repository to track and facilitate developments in this emerging field: https://github.com/Kali-Hac/AI4NDD-Survey..
keywords:
@@
Artificial intelligence, Neurodegenerative diseases, Gait, Parkinson’s disease, Alzheimer’s disease, Disease diagnosis
[…]
1 Introduction
@@
Neurodegenerative Diseases (NDs) such as Alzheimer’s disease (AD) and Parkinson’s disease (PD) are among the most widespread and devastating disorders affecting millions of people worldwide [1]. According to the Alzheimer’s Disease Association [2] and the Parkinson’s Foundation Report [3] in 2023, nearly 6.2 million people and 1 million people are diagnosed with AD and PD respectively in the United States. While in China, there are over 9.8 million AD patients [4], and it is estimated that Chinese PD patients will increase to 4.9 million by 2030, accounting for a half of the worldwide PD population [5]. From 1990 to 2019, the effected global population of AD, PD, and other neurodegenerative diseases ($e.g.$, amyotrophic lateral sclerosis) has significantly increased (illustrated in Fig. 2), leading to a growing burden on global healthcare system. On the one hand, these diseases could result in pathological gaits and chronic pain in body joints, tissues, and nerves [6, 7], which severely reduce the flexibility, stability, mobility, and other functional capabilities of a human body. On the other hand, they also impose large psychological and financial burdens to patients and their family [8, 9, 10]. The treatment cost for these diseases has reached approximately USD 130 billion per year and further rise has been estimated [11, 12]. Therefore, it is essential to conduct an earlier and more reliable diagnosis of these diseases, so as to provide a timely proper treatment to mitigate the burden of societal and healthcare resources.
@@
Figure 1: Worldwide effects (1990 versus 2019) of two most representative neurodegenerative diseases (AD and PD), measured in “Disability-Adjusted Life Years (DALYs)” [13].
@@
Figure 2: The number of studies (from 2012 to 2022) relevant to gait-based neurodegenerative disease diagnosis using AI.
@@
As one of the most essential motor symptoms associated with pathology, locomotion anomalies, especially observed in human gait, can reflect the incidence and progression of different NDs [14, 15, 16]. Taking AD, PD and amyotrophic lateral sclerosis (ALS) as an example, these diseases often cause patients to exhibit marked alterations and abnormalities in their walking patterns [17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. In general, compared with a normal gait (see Fig. 4), an impaired gait may display changes in speed, cadence, and limb moments within a gait cycle [27, 28, 19]. As shown in Fig. 4, the gait with a widened base, unsteadiness and irregularity of steps, and lateral veering (termed Parkinsonian gait) could suggest PD [19, 29]. In the case of advanced ALS, the patients may possess foot drop, where one foot flops down when lifting the leg, necessitating a higher lifting of the knee to prevent the toe from dragging on the ground (termed neuropathic gait) [30, 31, 32].
@@
In recent years, driven by the widespread availability of body positioning technologies and economical portable devices to monitor human gait, gait-based NDs diagnosis ($i.e.,$ NDs classification, identification or prediction) has propelled a surge of attention across academic, industrial, and medical communities [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45]. To perform NDs diagnosis from diverse gait data, the Artificial Intelligence (AI) technologies, especially Machine Learning (ML) and Deep Learning (DL) models, have been extensively explored and applied [46, 34, 35, 47, 36, 37]. A common practice is to extract gait parameters ($e.g.,$ stride length, stance duration, gait cycle time) from sensor-based data such as vertical ground reaction force (vGRF) or vision-based data such as gait images, and leverage ML algorithms ($e.g.,$ support vector machine (SVM)) or deep neural networks to automatically recognize gait patterns associated with NDs to aid in their diagnosis [48, 49, 50]. Compared with traditional diagnosis that requires manual observations of clinical motor symptoms ($e.g.,$ rigidity), these AI-based methods can analyze the abnormality in gait to automatically identify NDs related patterns, which shows immense potential in assisting healthcare professionals to diagnose and treat these diseases more efficiently [36, 37].
In this paper, we systematically review recent advancements of utilizing AI to help diagnose NDs from human gait. In particular, our work focuses on the diagnosis of five most prevalent and typical NDs (PD, AD, ALS, Huntington’s Disease (HD), Multiple Sclerosis (MS)) and all existing AI models that can help diagnose these diseases ($i.e.,$ identify, classify or predict NDs) from human gait data. A total of 2069 potentially relevant articles, including 14 survey papers, are found from three databases (PubMed, Web of Science (WoS), Google Scholar). We first screen these articles by titles and abstracts based on the pre-defined inclusion criteria, and then conduct a full-text assessment to exclude irrelevant papers. Finally, 164 of 2069 papers are selected to be included in our survey.
@@
Figure 3: Different phases of normal human gait with coordinated movements of legs and arms. Right: Examples for hemiplegic gait.
@@
Figure 4: Examples for hemiplegic gait (left), Parkinsonian gait (middle), and neuropathic gait (right) [51].
1.1 Objectives and Contributions
@@
The objectives of this survey can be encapsulated in the following six key questions (Q): (Q1) How is human gait associated with NDs? (Q2) What is the process of AI-assisted NDs diagnosis from gait data? (Q3) What are various gait data types, gait collecting approaches, and AI model types for NDs diagnosis? (Q4) What are state-of-the-art AI models and their performance results ($e.g.,$ diagnostic accuracy) in gait-based NDs diagnosis? (Q5) What are the current main challenges and promising future directions of this field? (Q6) Is it feasible to exploit 3D skeleton data to represent human gait and build AI models for NDs diagnosis?
@@ 2 Preliminaries
2.1 Basic Concepts
  1. 1.
    Gait: It is defined as a mechanism of locomotion, which contains rhythmic and coordinated movements of different limbs to enable the forward progression of the body [52, 53] (as shown in Fig. 4). The gait of a person typically consists of the motion of lower extremities and their correlated posture in upper limbs or torso, which normally requires a synthesized coordination of nerves, skeletons, and muscles [54]. Owing to this reason, the abnormality or impairment of gait can serve as an effective indicator of potential body pains or diseases caused by physical injury, aging or related disorders [55, 43].
  2. 2.
@@ 2.1 Basic Concepts
  5. 5.
    Disease Management: General disease management methods consists of therapies using a combination of medications, physical therapies, psychological therapies, interventional procedures or surgeries. For NDs, the management is often disease-specific with focus on either the disease pathogenesis or symptoms experienced [56].
2.2 Pathological Gaits
2.2.1 Neurodegenerative Disease Related Gaits
@@
2.2.2 Other Disease-Related Gaits
@@
Figure 5: Overview for the process of AI-assisted gait-based neurodegenerative disease diagnosis.
Figure 6: The taxonomy of gait data types used for AI-assisted NDs diagnosis.
Figure 7: Overview of AI model learning workflow.
2.3 Overview of AI-assisted Gait-Based Neurodegenerative Disease Diagnosis
@@
It is noteworthy that exploiting AI models for gait analyses and classification ($i.e.$, Step 2) is the main part of the gait-based NDs diagnosis process. It includes the workflow of designing, training, and validating AI models using pre-collected gait data. As shown in Fig. 7, the AI learning workflow can be divided into (a) raw gait data collection, (b) data preprocessing or/and augmentation, (c) model design or selection, (d) model training and validation, (e) model testing and deployment, and (f) iterative improvement ($e.g.,$ model fine-tuning using more data). For conventional ML models, feature engineering is usually required after step (c) to manually extract discriminative feature such as gait descriptors. For DL models, the design of models ($e.g.$, neural network architectures) (step (c)) and the sufficiency of training data (step (a) and (b)) are often the two most essential determinants of model performance. Since more high-quality training data can better guide the model optimization, extend the learnable feature space, and improve the generalization ability ($i.e.,$ higher performance on new testing data) of the model, it is feasible to collect more gait data while using effective data augmentation strategies ($i.e.$, generate more augmented training samples) to iteratively boost the model performance (detailed in Sec. 5).
2.4 Gait Data Taxonomy
@@
  • Sensor Modality: Advancement in sensor technologies allows its application in various tasks such as gait analysis, action recognition, etc. The sensors explored among the collected papers offer various advantages and could be classified into these four types, including force, inertial, hybrid, and other less prevalent sensors [57, 58, 59, 60, 41].
    1. 1.
@@
      Force Sensor (FS): There are mainly two types of force or pressure capturing approaches. The first is to use non-wearable floor-mounted devices ($e.g$, force mats, treadmill) to measure vertical Ground Reaction Force (vGRF), moments, or plantar pressure profiles [41]. In a laboratory setting, these stationary floor-mounted devices are sufficient for precise data collection. However, since those devices are often bulky, operating-costly, and expensive, it’s not often clinically applicable. Therefore, researchers develop some small wearable devices ($e.g.$, insole) whose application in the field has became a popular choice and risen significantly in the past decade, due to its portable features and cheap installment, at the cost of less accuracy and reliability compare with the floor-mounted ones [61, 62].
    2. 2.
@@
      Inertial Sensor: Also termed as Inertial Measurement Unit (IMU), is often attached to limb segments ($e.g$, foot) [63] or embedded in wearable devices ($e.g.$, smartwatch) to capture dynamic motion ($e.g.$, 3D-acceleration) of the subject [64, 65]. These sensors or devices record the repeating pattern of gait signals generated by the locomotion of the subjects. These signals will then be interpreted with various algorithms for the intended tasks, including gait event detection, kinetic or kinematic parameters estimation, and gait classification.
    3. 3.
@@
      Multi-Sensor: These sensors typically combine various types of sensors for a more comprehensive data collection. Among the surveyed papers, studies such as [66, 67, 68] utilize both IMU sensors and Force Sensitive Resistors (FSRs), while [69] employs Electroencephalogram (EEG) sensors and FSRs. These combinations harvest the strengths of each sensor type hence help enhancing the reliability and accuracy of gait event detection and classification.
    4. 4.
@@
      Other Sensors: Sensors included in this category are less prevalent and exploited among collected papers. However, these sensors still play an important role in providing a more full-sided understanding of the human gait. For instance, EEG sensor is used to measure the brain electrical activity signals data which can be used to analyze brain activity patterns that may correlate with specific gait abnormalities [70, 69, 71]. Electromyography (EMG) sensor is another useful sensor which is used to measure the electrical activity produced by muscles whose activities are controlled by the motor neurons [72]. Since NDs patients suffer from the degeneration of nerve cells, they manifest significantly different muscle electrical pattern from healthy subjects [36]. Therefore, their EMG data can be utilized to aid NDs diagnosis [72].
@@
  • Vision Modality: The vision modality employs optoelectronic Motion capture (Mocap)systems, including various types of cameras ($e.g.$, Infrared (IR) Camera [73]), to capture the locomotion of subjects during walking. It’s used to overcome the limitations of the traditional clinical examination which purely depends on the physicians’ naked eyes. Depends on the precision requirement of the video data, the vision modality could be further divided into two sub-categories, namely marker-based and marker-less systems [47].
    1. 1.
@@
      Marker-based: This approach is developed at an early stage when the cameras are not advanced enough for precise motion capture. Therefore, these Mocap systems ($e.g.$, Vicon) use reflective markers to help capture targets’ movement in the images more precisely. The advantages of using the marker-based approach are that they’re highly accurate and provides exact location of the body landmark. These advantages hence make this modality a gold standard in the field. However, this method requires attaching markers to the body, leading to additional expenses and the potential of alter the gait [74, 75, 76].
    2. 2.
@@
  • Sensor+Vision Modality: This modality could also be referred to as multi-modal modality. In comparison with single-modal approaches, although the latter ones have their own merits and show some promising results, this combined approach unleashes a bigger potential, such as providing richer data and more comprehensive features, capturing more subtle changes in gait, improving recognition performance [71, 77, 78]. In particular, the application of Microsoft Kinect system has gain a tremendous amount of attention in the field due to its utilization of both depth sensor and RGB camera. This integration enables a more accurate estimation of the gait parameters without sacrificing the advantages mentioned in the marker-less vision modality. More specifically, using conventional cameras to estimate skeleton coordinates produces coarse results. However, when employing the depth sensor, the spatiotemporal features can be precisely derived from the skeleton joints’ coordinates obtained using it. The precision and efficacy of using Kinect has been verified in many studies which compare the results with the gold standard (marker-based vision) approach. Nevertheless, one limitation of current Kinect system is that its precision reduces as the subject walks further away from the camera, resulting in a width and length-limited data collection space [79, 80, 81, 82]. Besides the RGB-D camera approach, there are also other combinations of these two modalities. For example, Tahir et al. fused a marker-based IR camera and a force-sensitive platform [83], while Chatzaki et al. utilized a conventional camera, IMU, and a force platform [78]. Additionally, Zhao et al. merge some vGRF and IMU sensors with Kinect [77].
  • Other Modalities: Trabassi et al. use Electronic Medical Record (EMR) which contains various patient-related features and medical history data that are indicative of early PD symptoms or risk factors as well as the first diagnosis of gait or tremor disorders to create prediction models [84].
@@
Figure 8: Statistics of the included papers with different (a) AI model types, (b) NDs types, (c) experimental sample sizes, (d) gait data types, (e) sensor types, and (f) countries. We show the lowest and top three quantities in (f). Papers of different years are shown in Fig. 2. Detailed analyses are provided in Sec. 3.2.
2.5 AI Model Taxonomy